OpenStudy (anonymous):

Can somebody please explain this to me?. How do you evaluate a^n if both a and n are positive integers? Does multiplying a by itself n times make sense if n is a negative integer? What do we do when n is negative? Why?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

hes very good.

jimthompson5910 (jim_thompson5910):

" How do you evaluate a^n if both a and n are positive integers?" you just multiply 'a' by itself n times example 7^3 = 7*7*7 = 343 we multiply 7 by itself 3 times

jimthompson5910 (jim_thompson5910):

it doesn't make sense if n is negative, but you can easily fix that using the rule \[\Large x^{-k} = \frac{1}{x^k}\] example \[\Large 7^{-3} = \frac{1}{7^3}\]